A college has $60,000 to invest in an endowment in order to have an annual income of $5000 for a scholarship. Part of this will be invested in government bonds at 8% and the remainder in long-term fixed deposits at 10.5%. How much should be invested in each to provide the required income?
Answer by checkley71(8403)

x*.08+(60000-x).105=5000 or .08x+6300-.105x=5000 or -.025x=-1300 or x=1300/.025 or x=5200 invested @ 8% & 60000-52000=8000 invested @ 10.5%